OpenStudy (anonymous):

the complex fifth roots of 5

OpenStudy (anonymous):

D:

OpenStudy (anonymous):

we'll need serious digging for this :3

OpenStudy (anonymous):

all complex numbers can be written in polar form. so how about 5?

OpenStudy (anonymous):

*snap* *snap* ... you there @shoyos8 ?

OpenStudy (anonymous):

Can you write as 5 = 5(cos 0 + i sin 0) ?

OpenStudy (science0229):

Right Now, @shoyos8 do you know the complex root formula?

OpenStudy (science0229):

Looks something like this:\[x^n=c=r(\cos \theta+i \sin \theta) \rightarrow x=r^{\frac{ 1 }{ n }}(\cos{\frac{ \theta+2k \pi }{ n }}+isin{\frac{ \theta+2k \pi }{ n }})\]

OpenStudy (science0229):

where k=0,1,...,n-1

OpenStudy (science0229):

c is constant

OpenStudy (science0229):

@shoyos8